# MambaFlow: A Mamba-Centric Architecture for End-to-End Optical Flow Estimation

### v1

@@ Abstract
Optical flow estimation based on deep learning, particularly the recently proposed top-performing methods that incorporate the Transformer, has demonstrated impressive performance, due to the Transformer’s powerful global modeling capabilities. However, the quadratic computational complexity of attention mechanism in the Transformers results in time-consuming training and inference. To alleviate these issues, we propose a novel MambaFlow framework that leverages the high accuracy and efficiency of Mamba architecture to capture features with local correlation while preserving its global information, achieving remarkable performance. To the best of our knowledge, the proposed method is the first Mamba-centric architecture for end-to-end optical flow estimation. It comprises two primary contributed components, both of which are Mamba-centric: a feature enhancement Mamba (FEM) module designed to optimize feature representation quality and a flow propagation Mamba (FPM) module engineered to address occlusion issues by facilitate effective flow information dissemination. Extensive experiments demonstrate that our approach achieves state-of-the-art results, despite encountering occluded regions. On the Sintel benchmark, MambaFlow achieves an EPE all of 1.60, surpassing the leading 1.74 of GMFlow. Additionally, MambaFlow significantly improves inference speed with a runtime of 0.113 seconds, making it 18% faster than GMFlow. The source code will be made publicly available upon acceptance of the paper.
@@
1 Introduction
Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence. Optical flow estimation holds significant and widespread applications in areas such as action recognition [42, 37], video interpolation [11, 34, 19], and autonomous driving [38].
Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow estimation. FlowNet [4], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm. However, it faced limitations in handling large displacements and complex motions, along with high computational costs. Subsequently, PWC-Net [30] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features. Nevertheless, it still encountered challenges in large displacement scenarios. To address these issues further, RAFT [31] introduced a recurrent iterative refinement mechanism, significantly enhancing the accuracy of optical flow estimation. Still, its optimization process relies on numerous iterations, leading to increased parameter counts and substantially longer inference time.
In recent years, the Transformer architecture, with its powerful global modeling capabilities, has been introduced into optical flow estimation tasks. For instance, GMFlow [36] and FlowFormer [10] leverage attention mechanisms to capture long-range dependencies, significantly improving the accuracy of optical flow estimation. However, the quadratic computational complexity of the Transformers significantly increases the inference time in high-resolution scenarios, severely limiting their efficiency in practical applications. This raises a critical question: How can we achieve efficient optical flow estimation with reduced computational overhead while preserving high accuracy?
Mamba [5], rooted in state space models (SSMs)[6], employs an efficient selective state space mechanism to excel in tasks requiring the integration of global information and handling complex spatiotemporal dependencies. Simultaneously, it achieves linear computational complexity when capturing long-range dependencies in sequences. Inspired, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges of large displacements, occlusions, and computational efficiency. As demonstrated in Fig. 1, our method significantly reduces inference time while maintaining high accuracy. The framework introduces two core innovations: a feature enhancement Mamba (FEM) module, which optimizes feature representations to enhance global matching performance, and a flow propagation Mamba (FPM) module, which propagates flow information to tackle challenges in occluded regions.
Specifically, our framework begins with a CNN backbone to extract robust features from the input image pair. We then propose the FEM module, which employs the Self-Mamba and Cross-Mamba mechanisms to strengthen the correlations between the extracted features, allowing for the capture of complex global information and long-range dependencies. Building on the enhanced features provided, a global matching module compute similarities and perform global optimization, to generate initial flow. This process fully utilizes global information, effectively tackling the challenges posed by large displacements. At last, we propose a Mamba-centric flow propagation method, the FPM module, which leverages feature similarity to guide propagation and produce the final accurate flow [36], tackling the occlusion issue in predicting optical flow.
@@ 1 Introduction
  • We propose MambaFlow, a novel optical flow estimation architecture built on the Mamba framework. To the best of our knowledge, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation. Our method achieves comparable accuracy to state-of-the-art algorithms while significantly reducing inference time. With a runtime of 0.113 seconds, our method is 18% faster than GMFlow and 10 times faster than MS-RAFT+.
  • We introduce a feature enhancement Mamba (FEM) module, which integrates Self-Mamba and Cross-Mamba mechanisms to address insufficient feature representation. This module established a new Mamba-centric paradigm for accurate and efficient feature enhancement.
  • We design a flow propagation Mamba (FPM) module for flow propagation that utilizes feature self-similarity to guide the final flow estimation. This module tackles the poor performance in recurrent refined methods by effectively propagating flow information, offering significant potential to enhance accuracy and speed, both currently and in the foreseeable future.
@@ 2.1 Optical Flow Estimation
The optical flow problem is traditionally approached as an energy minimization task. Based on the assumptions of brightness constancy and spatial smoothness, Horn and Schunck [8] introduced a variational method to tackle this problem. The following studies refined optical flow estimation by framing it as an optimization problem, improving accuracy and robustness through regularization that maximizes the visual similarity between corresponding pixels in the image. In the era of deep learning, following the introduction of FlowNet [4] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [30] and SpyNet [28], adopted a coarse-to-fine strategy for optical flow estimation. However, small displacements of fast movement were often missed via the inaccurate flow guidance at the coarse level. To overcome this obstacle, RAFT [31] proposed an iterative refinement method using full-field, achieving significant improvements. Constructing upon the innovation of RAFT, later works such as GMA [15] improved the handling of occluded regions by introducing a global motion aggregation module. Similarly, SeparableFlow [40] introduced a separable cost volume module that used a global motion aggregation layer to utilize global contextual cues and prior knowledge to resolve motion ambiguity in regions. Despite the advancements, CNN-based models, including those built upon the RAFT framework, still rely on multiple iterations of refinement to achieve high-precision optical flow estimation. This iterative process significantly increases computational overhead and slows inference speed, thereby limiting their practicality for efficiency-sensitive applications.
In recent years, the Transformer architecture has demonstrated exceptional proficiency in modeling long-range dependencies and global information through its self-attention mechanism, making it a popular choice for optical flow estimation tasks. By incorporating more global and contextual information, it has significantly improved estimation accuracy. FlowFormer [10] further optimized model performance by encoding 4D cost tags into cost memory and applying alternating groups of Transformer layers in the latent space. CRAFT [29] strengthened the globality and semantic stability of features by leveraging semantic smoothing Transformer layers. GAFlow [22] added Gaussian attention to the existing Transformer block to emphasize local attributes during representation learning, highlight local neighborhoods containing fine-grained structural information, and strengthen motion affinity during matching. In addition, many models took advantage of the Transformer’s global modeling capabilities, adopting a global matching strategy to avoid multiple iterations of refinement, significantly enhancing efficiency. GMFlow [36] reformulated the optical flow problem as a global matching problem and obtained the optical flow using one-time global matching, which was more accurate and faster. Transflow [21] used spatial self-attention and cross-attention mechanisms between adjacent frames to effectively captured global dependencies, providing more accurate correlations and reliable matches in flow estimation. Still, despite these advances in accuracy, the Transformer-based optical flow methods face challenges due to the quadratic complexity of the attention mechanism led to huge computational complexity and memory requirements. This slows inference speed and restricts their use in resource-limited settings, creating practical bottlenecks.
In our proposed MambaFlow architecture, we employ the Mamba module, which contains linear complexity and discriminative feature representation. This approach significantly reduces the overhead of computational resources and memory requirements while maintaining high accuracy comparable to Transformer-based methods, A remarkable breakthrough for the practical application of optical flow models.
2.2 State Space Model in Optical Flow
@@
State space models [7], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies. Mamba [5] introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language. Related work on Mamba [33] has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity. This superiority led to further exploration like Vision Mamba [44] and Vmamba [20], which applied SSMs to visual tasks such as image classification [20, 44], video understanding [17, 32], biomedical image segmentation [23, 35] and others [9, 12, 26, 27, 39]. However, their applicability to optical flow estimation remained unclear until recently. Lin et al. [18] pioneered Mamba into the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder. This approach can effectively integrate global motion information in the point cloud and significantly improved inference speed.
Inspired by the Mamba’s linear complexity, we propose a Mamba-centric optical flow estimation architecture, aiming to replace the traditional attention mechanism and thus improve computational efficiency of the model. In this work, we explore the potential of Mamba for modeling sequential interactions in optical flow tasks and propose FEM module and FPM module. The FEM module enhances the features from pairs of images, capturing the dependency between them, while the FPM module refines optical flow predictions for occluded pixels by leveraging both the optical flow field and image structure similarity. By utilizing Mamba’s linear time complexity, our approach overcomes the quadratic complexity limitations of traditional Transformer models when processing long sequences, resulting in significantly improved inference efficiency.
Figure 2: The overall network architecture of our MambaFlow. First, we perform downsampling using convolutional layers and input the initially extracted features into the feature enhancement Mamba group (FEMG) module, which is composed of multiple feature enhancement Mamba (FEM) module. Each FEM module consists of Self-Mamba, Cross-Mamba, and MLP for feature enhancement. Next, global matching is applied to the output feature matrix for similarity matching, and the original optical flow is obtained through a softmax layer. Finally, the flow propagation Mamba (FPM) module is used to address inaccuracies in optical flow prediction caused by occlusion.
@@ 3.1 Preliminaries
Figure 3: The architecture of feature enhancement Mamba (FEM) module. It consists of three core components: Self-Mamba (a), which focuses on a single feature input, Cross-Mamba (b), which handles the interaction between two features, and MLP (c), which performs non-linear feature transformation.
@@ 3.2 Overall Architecture
Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between images. As shown in Fig. 2, our MambaFlow consists of three components: feature enhancement Mamba, global matching and flow propagation Mamba.
Given two consecutive frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we apply a shared-weight convolutional neural network to perform downsampling, obtaining dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$. Here, $H,W,$ and $D$ denote the height, width, and feature dimension of the feature maps, respectively. Next, the image features are processed through the feature enhancement stage to obtain the enhanced features of the $l$-th layer, $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}\in\mathbb{R}^{H\times W\times D}$. This stage is composed of multiple stacked FEM modules, each consisting of Self-Mamba and Cross-Mamba blocks. Additionally, a Multi-Layer Perceptron (MLP) layer is introduced at the end of each group to further integrate local and global features and enhance feature representation. Finally, we employ the Global Matching method to predict optical flow and use our FPM module to significantly improve performance in occluded regions.
@@ 3.3 Self-Mamba
The input features $\boldsymbol{\mathit{F}}_{1},\boldsymbol{\mathit{F}}_{2}\in\mathbb{R}^{H\times W\times D}$ are enriched with learnable positional embeddings to retain spatial positional information. Subsequently, $\boldsymbol{\mathit{F}}_{1}$ and $\boldsymbol{\mathit{F}}_{2}$ are processed through a weight-sharing Self-Mamba block, generating $x_{q}$ and $x_{v}$, respectively.
As shown in Fig. 3 (a), the processing within each Self-Mamba block can be summarized as a bidirectional operation, where the forward and backward hidden states are computed and combined to enrich the feature representation. The overall process is expressed as:
@@
Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $x$ being reversed to distinguish between the forward and backward directions.
@@ 3.4 Cross-Mamba
Mamba is specifically designed to capture long-term dependencies within a single sequence, but it cannot model the interrelations between two sets as effectively as the cross-attention mechanism in the Transformers. To overcome this limitation, we propose Cross-Mamba, an extension of the Mamba framework that adapts the cross-attention mechanism to better capture the interactions between two distinct sets. This idea, which builds upon the work by [33], is further extended in our approach. Since the parameters $\overline{\mathbf{A}}$, $\overline{\mathbf{B}}$, ${\mathbf{C}}$, and $\boldsymbol{\Delta}$ are associated with the input, $\overline{\mathbf{K}}$ in (5) is extended into an $M\times M$ matrix [1]. The element $\alpha_{i,j}$, located at row $i$ and column $j$, is determined as follows:
$\alpha_{i,j}=\mathbf{C_{i}}\prod_{k=j+1}^{i}\overline{\mathbf{A}}_{k}\overline{\mathbf{B}}_{j}.$ (11)
Substituting the equations from (3) to (9), we can ultimately derive the output of Mamba as follows:
$y_{i}=\sum_{j=1}^{i}\alpha_{i,j}x_{j}=\sum_{j=1}^{i}Q_{i}K_{j}H_{i,j}V_{j},$ (12)
where
$Q_{i}$ $=S_{C}(x_{i}),$ (13)
$\begin{split}K_{j}=&\left(\text{Softplus}(S_{\Delta}(x_{j}))\mathbf{A}\right)^{-1}\\ &\left(\exp\left(\text{Softplus}(S_{\Delta}(x_{j}))\mathbf{A}\right)-\mathbf{I}\right)\\ &\left(\text{Softplus}(S_{\Delta}(x_{j}))S_{B}(x_{j})\right),\end{split}$ (14)
$H_{i,j}$ $=\prod_{k=j+1}^{i}\exp\left(\text{Softplus}(S_{\Delta}(x_{k}))\mathbf{A}\right),$ (15)
$V_{i}$ $=x_{i}.$ (16)
Cross-attention is a crucial mechanism for modeling the relationships between two sets. It operates by computing interactions between the query ($Q$) from one set and the keys ($K$) and values ($V$) from another set, thereby identifying and focusing on the most relevant elements in the target set.By applying $x_{q}$ to (13), and $x_{v}$ to (14), (15) and (16), respectively, Cross-Mamba can be expressed in the following form:
$y_{i}=\sum_{j=1}^{i}f_{m,q}(x_{q,i})f_{m,k}(x_{v,j})H_{i,j}x_{v,j}.$ (17)
Cross-Mamba extends the original Mamba by expanding its application from a single input sequence to two input sequences, as shown in Fig. 3 (b). It is formally defined as follows:
$y=\text{Cross-Mamba}(x_{q},x_{v}).$ (18)
Similarly, we introduce the bidirectional mechanism of Self-Mamba into Cross-Mamba. The formulation is expressed as follows:
[…]

### v2

@@ Abstract
Recently, the Mamba architecture has demonstrated significant successes in various computer vision tasks, such as classification and segmentation. However, its application to optical flow estimation remains unexplored. In this paper, we introduce MambaFlow, a novel framework designed to leverage the high accuracy and efficiency of the Mamba architecture for capturing locally correlated features while preserving global information in end-to-end optical flow estimation. To our knowledge, MambaFlow is the first architecture centered around the Mamba design tailored specifically for optical flow estimation. It comprises two key components: (1) PolyMamba, which optimizes feature representation; and (2) PulseMamba, which facilitates efficient flow information dissemination. Our extensive experiments demonstrate that MambaFlow achieves remarkable results. On the Sintel benchmark, MambaFlow records an endpoint error (EPE) of 1.43 and an inference speed of 0.113 seconds, surpassing the state-of-the-art methods including GMFlow (with 18.9% lower EPE and 18.1% faster inference), SeparableFlow (5% lower EPE and 50.5% faster), CRAFT (1.11% lower EPE and 76.5% faster), and DIP (0.7% lower EPE and 77.2% faster)—demonstrating stronger potential for real-world deployment on resource-constrained devices. The source code will be made publicly available upon acceptance of the paper.
1 Introduction
Optical flow estimation, a fundamental problem in computer vision, aims to compute the motion vectors of pixels between consecutive frames in a video sequence. It holds significant and widespread applications in areas such as action recognition [65, 56], video interpolation [16, 51, 28], autonomous driving [59], etc.
Architectures based on Convolutional Neural Networks (CNNs) have achieved pioneering progress in the field of optical flow. FlowNet [5], as the first end-to-end CNNs-based flow network, revealed the immense potential of this learning paradigm. However, it faced limitations in handling large displacements and complex motions. Subsequently, PWC-Net [44] improved estimation accuracy in complex motion scenarios by incorporating pyramid feature extraction and a cost-volume construction, which effectively captured multi-scale local features. Nevertheless, it still encountered challenges in large displacement scenarios. To address these issues, RAFT [45] introduced a recurrent iterative refinement mechanism, which significantly enhanced the accuracy of optical flow estimation. However, such innovation brought a new challenge, which its optimization process relies on numerous iterations, leading to increased parameter counts and longer inference time. Many subsequent variants of the RAFT algorithm have attempted upgrates, such as MS-RAFT [19] and SEA-RAFT [49], but their ability to handle large displacements and complex motions still requires further improvement. To tackle the aforementioned issues, the Transformer architecture, known for its powerful global modeling capabilities, has been introduced into optical flow estimation tasks [15, 42, 30, 43]. Most of them leverage attention mechanism to enhance feature representation, which capture long-range dependencies to significantly improve the accuracy. Notably, unlike optical flow algorithms based on RNNs for flow decoding [45, 20], GMFlow [54] utilizes the attention mechanism to iteratively decode the final optical flow, resulting in faster flow refinement. However, they encounter the quadratic computational complexity of Transformers both during the training and inference stages [54, 43, 15, 42, 30], which severely limits their efficiency in practical applications.
@@ 1 Introduction
Recently, Mamba [8, 10] has been widely used in various computer vision tasks, such as image classification [7, 57, 58],image segmentation [27, 67, 6], object detection [33, 32, 46], image generation [13, 23], video understanding [21, 24, 52], etc. Those remarkable achievements have demonstrated that Mamba can effectively address the challenges posed by CNNs and Transformers, including the local perception limitations of CNNs and the quadratic computational complexity of Transformers [55, 63]. It benefits from a selective state space mechanism, which can efficiently capture global information and handle complex spatiotemporal dependencies with linear complexity [63]. Specifically, Mamba has been conceptually summarized as being ideally suited for tasks involving long sequences and autoregression [60] due to its inherent RNN mechanism within the SSM [10, 9, 8].
Inspired by Mamba’s capabilities for long sequence modeling and autoregression with linear complexity, we propose a novel optical flow estimation framework, MambaFlow, designed to address the challenges in CNNs or Transformer-based methods. As demonstrated in Fig. 1, our method significantly reduces inference time while maintaining remarkblely high accuracy.The framework introduces two core components: (1) the PolyMamba module, which utilizes Self-Mamba, Cross-Mamba, and Multi-Layer Perceptron (MLP) to globally optimize feature representations with lightweight computation, and (2) the PulseMamba module, which effectively combines motion features with Mamba’s autoregressive ability to iteratively generate flow information, addressing challenges such as large displacement and occluded regions.
Specifically, we begins with a CNN backbone to extract robust features from the input image pairs similar to traditional practice [49, 54, 20]. Subsequently, we leverage the suggested PolyMamba to enhance the extracted features, allowing for the capture of complex global information and long-range dependencies. Building on the enhanced features, we further utilize a global matching mechanism [54, 64] to compute similarities for initial flow. At last, we employ PulseMamba to recurrently refine and yield the final precise flow [54]. Our algorithm uses Mamba as the core components in both the encoding and decoding stage, making it an end-to-end optical flow estimation algorithm centered around Mamba. By leveraging Mamba’s capabilities in global modeling and autoregressive properties, as well as its computationally efficient characteristics, the proposed approach achieves a good trade-off between accuracy and performance.
@@
  • We propose MambaFlow, a novel optical flow estimation architecture built on the Mamba framework both in feature enhancement and flow propagation. To our knowledge, this is the first Mamba-centric architecture designed for end-to-end optical flow estimation. Our method achieves accuracy comparable to state-of-the-art algorithms while significantly reducing training and inference time, which makes it more suitable for real-world application on resource-constrained devices.
  • We introduce the PolyMamba module, which consists of a Self-Mamba, a Cross-Mamba, and a MLP in sequence, to globally enhance feature representation for both the reference image and its cross features with the matching image. As we will demonstrate, it effectively leverages the Mamba’s long sequence modeling capabilities while maintaining a linear computational load.
  • We design the PulseMamba, which includes a motion encoder and a Mamba to facilitate flow propagation, addressing the computational overload, large displacement and occulusion issues in recurrently refined RNN-like methods. By leveraging the Mamba’s powerful autoregressive capability and linear complexity, it offers significant potential to enhance accuracy and speed simultaneously through ablation studies.
@@ 2.1 Optical Flow Estimation
Optical flow estimation was traditionally formulated as an energy minimization task [1, 12, 38], with early variational methods like Horn and Schunck [12] relying on brightness constancy and spatial smoothness assumptions. Subsequent optimization-based approaches improved robustness through regularization [2].
In the era of deep learning, following the introduction of FlowNet [5] as the first CNN-based optical flow method, numerous subsequent algorithms such as PWC-Net [44] and SpyNet [40], adopted a coarse-to-fine strategy for optical flow estimation. However, these approaches struggled with fast motions and small displacements due to inaccurate coarse-level guidance. RAFT [45] addressed this via iterative full-field refinement, inspiring variants such as GMA [20], SeparableFlow [62],and MS-RAFT [19], which introduced modules to better handle occlusion and motion ambiguity. Despite their accuracy, these CNN-based methods require heavy iterative computation and struggle with large displacements and complex motion, limiting their efficiency on resource-constrained devices.
Transformers have emerged as powerful alternatives due to their ability to model global dependencies via self-attention. Models like FlowFormer [15], CRAFT [43], and GAFlow [31] leveraged Transformer architectures to enhance feature expressiveness and global matching. GMFlow [54] and TransFlow [30] further improved efficiency by reformulating optical flow as a one-shot global matching problem. However, the quadratic complexity of attention mechanisms results in high computational and memory costs, limiting deployment on lightweight systems.
2.2 State Space Models
@@
State space models [11], which originated from control theory, have been effectively integrated with deep learning to capture long-range dependencies. Mamba [8] introduced an input adaptation mechanism to enhance the state space model, providing faster inference speed, higher throughput and overall indicators compared to the Transformers of the same size, especially in natural language. Related work on Mamba [50] has shown that it has comparable performance to the Transformer-based methods while significantly reducing computational complexity. This superiority led to further exploration like Vision Mamba [68] and Vmamba [29], which applied SSMs to visual tasks such as image classification [29, 68], video understanding [25, 47], biomedical image segmentation [34, 53] and others [14, 17, 37, 39, 61]. However, their applicability to optical flow estimation has remained unexplored until recently. Lin et al. [26] pioneered the use of Mamba in the scene flow estimation task, proposing an SSM-based iterative update module to replace the GRU module as a decoder. Nevertheless, it was specifically designed for scene flow with point cloud feature and is not suitable for the generally used backbone optical flow methods.
Inspired by the Mamba’s strength of long sequence modeling and autoregression with linear complexity, we propose a Mamba-centric architecture for optical flow estimation. Our goal is to design an end-to-end solution that utilizes Mamba for feature enhancement and flow optimization. During the feature enhancement phase, PolyMamba enhances features from pairs of images by sequentially employing Self-Mamba, Cross-Mamba, and MLP to capture global information and mutual dependencies. At the flow refinement phase, PulseMamba iteratively optimize the fina optical flow by leveraging a motion encoder and a SSM. These Mamba-centric designs guarantee the accuracy of our method while maintaining its high speed in training and inference, providing valuable insights for resource-constrained devices.
Figure 2: The overall network architecture of our MambaFlow.
@@ 3.2 Overall Architecture
Optical flow can be intuitively understood as a task aimed at finding pixel-wise correspondences between consecutive images. As shown in Fig. 2, our proposed framework, MambaFlow, is centered around two novel modules: the feature enhancement module PolyMamba and the flow refinement module PulseMamba.
Given two input frames $\boldsymbol{I}_{1}$ and $\boldsymbol{I}_{2}$, we first extract dense feature maps $\boldsymbol{F}_{1},\boldsymbol{F}_{2}\in\mathbb{R}^{H\times W\times D}$ using a shared-weight convolutional encoder, where $H$, $W$, and $D$ denote the height, width, and feature dimension of the features. These features are then processed by the PolyMamba module, which stacks Self-Mamba and Cross-Mamba blocks to capture both intra-frame and inter-frame dependencies. To further enhance representation capacity, a lightweight MLP layer is appended after each block group for local-global feature fusion. The resulting features $\boldsymbol{\mathit{F}}_{q}$ and $\boldsymbol{\mathit{F}}_{v}\in\mathbb{R}^{H\times W\times D}$ are then used for global matching to estimate the initial flow.
The initial prediction $\mathbf{V}_{\text{initial}}$ is then refined by the PulseMamba module, which incorporates cost volume and motion features into a stack of Mamba layers for iterative flow refinement. This module significantly improves performance in occluded or textureless regions, enhancing both the accuracy and robustness of optical flow estimation.
@@ 3.3 Self-Mamba
The input features $\boldsymbol{\mathit{F}}_{1},\boldsymbol{\mathit{F}}_{2}\in\mathbb{R}^{H\times W\times D}$ are enriched with learnable positional embeddings to retain spatial positional information. Subsequently, $\boldsymbol{\mathit{F}}_{1}$ and $\boldsymbol{\mathit{F}}_{2}$ are processed through a weight-sharing Self-Mamba block, yielding enhanced representations denoted as $\boldsymbol{\mathit{F}}_{1}^{\text{self}}$ and $\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, respectively.
The processing within each Self-Mamba block can be summarized as a bidirectional operation, where the forward and backward hidden states are computed and combined to enrich the feature representation. The overall process is expressed as:
@@
Here, the Self-Mamba function represents the state space model used to compute the forward and backward hidden states, with the sequence $\boldsymbol{\mathit{F}_{i}}$ being reversed to distinguish between the forward and backward directions.
@@ 3.4 Cross-Mamba
While the Self-Mamba architecture is effective in modeling long-range dependencies within a single sequence, it lacks the capacity to capture inter-sequence relationships—an ability that is crucial in optical flow estimation, where information from two correlated feature maps must be jointly reasoned. To address this limitation, we introduce a cross-sequence interaction module based on selective state-space modeling, enabling rich bidirectional information exchange between paired features.
Given the enhanced features $\boldsymbol{\mathit{F}}_{1}^{\text{self}},\boldsymbol{\mathit{F}}_{2}^{\text{self}}$, the main branch applies a depth-wise causal convolution and activation:
$\boldsymbol{x}_{\mathrm{conv}}=\texttt{SiLU}\left(\texttt{DWConv}(\boldsymbol{\mathit{F}}_{1}^{\text{self}})\right),$ (11)
while the modulation branch linearly projects the second stream:
$\boldsymbol{x}_{\mathrm{mod}}=\texttt{Proj}(\boldsymbol{\mathit{F}}_{2}^{\text{self}}).$ (12)
Their concatenation is projected into three parameter groups:
$[\,\boldsymbol{\Delta},\;\mathbf{B},\;\mathbf{C}\,]=\texttt{Linear}([\boldsymbol{x}_{\mathrm{conv}};\;\boldsymbol{x}_{\mathrm{mod}}]),$ (13)
The hidden state $\boldsymbol{y}$ is then updated through a SelectiveScan kernel, which dynamically integrates temporal and spatial information using parameterized state-space operations.
A final linear projection yields the cross-attended representation:
$\boldsymbol{\mathit{F}}^{\text{cross}}=\texttt{OutProj}(\boldsymbol{y}).$ (14)
To model temporal symmetry, Cross-Mamba is applied in both forward and backward directions—just as in the Self-Mamba block—and the outputs are summed: